OpenStudy (anonymous):

Use the elimination method to solve the system of equations. a+6b=5 a+3b=-1

OpenStudy (anonymous):

You would multiply the bottom equation by -1 to get rid of the "a's". Then add the two equations together and solve for b:)

OpenStudy (anonymous):

So! a+6b=5 -a-3b=1 That would cancel out the "a's":)

OpenStudy (anonymous):

so a=-7 and b=2??

OpenStudy (anonymous):

Yessmam:) Hope I helped:)

OpenStudy (anonymous):

Yes, thanks